# Fight like a Woman:

### v1

@@ Abstract
This article investigates the impact of judges’ gender on the outcome of domestic violence cases. Using data From São Paulo, Brazil, between 2010 and 2019, we compare conviction rates by judge’s gender and find that a domestic violence case assigned to a female judge is 31% (10 p.p) more likely to result in conviction than a case assigned to a male judge with similar career characteristics. To show that this decision gap rises due to different gender perspectives about domestic violence instead of rising due to female judges being tougher than male judges, we compare it against gender conviction rate gaps in similar types of crimes. We find that the gender conviction rate gap for domestic violence cases is significantly larger than the same gap for other misdemeanor cases (3 p.p. larger) and for other physical assault cases (8 p.p. larger). Lastly, we find evidence that at least two channels explain this gender conviction rate gap for domestic violence cases: gender-based differences in evidence interpretation and gender-based sentencing criteria.
@@ Introduction
Violence against women is frequently perpetrated inside the home by the ones closest to them. The United Nations Office on Drugs and Crime (2023) estimates that 55% of all homicides against women worldwide in 2022 were committed by intimate partners or other family members. As this violence is a criminal offense, a key component of the fight against domestic violence is the Criminal Court System and, therefore, how court judges approach this type of crime (Beecher-Monas, 2001). For instance, the HumanRightsWatch (1991) argues that, in the 1980s and early 1990s, Brazilian judges failed to analyze crimes of domestic violence against women in a non-discriminatory manner. [^1] In the years since, Brazilian authorities have made important legal and institutional changes to address domestic violence crimes. The most prominent of these is the Maria da Penha Law, enacted in 2006, which characterizes domestic violence as a specific type of criminal offense. Moreover, Kafka et al. (2019) finds that American judges in North Carolina use gender stereotypes to argue that some women want to “game the system” with “frivolous cases” when requesting protective orders against their abusive partners.
@@
In this analysis, we find that a domestic violence case assigned to a female judge is 31% (10 p.p) more likely to result in conviction than a case assigned to a male judge with similar career characteristics. Since this decision gap may be due to differences between male and female judges regardless of their perspectives on domestic violence, we compare this gap to gender decision gaps for two other types of crime: misdemeanors and physical assault. Because we focus on domestic violence crimes that do not result in death or long-term physical impairment, misdemeanors are comparable to domestic violence cases in their sentence severity (according to Brazilian Law). Physical assault cases are comparable to domestic violence cases because both involve some sort of physical aggression. We find that the gender conviction rate gap for domestic violence cases is more than twice the size of that of other misdemeanors. There are no differences between female and male judges in sentencing physical assault cases that do not involve domestic violence. Consequently, the gender conviction rate gap for domestic violence cases likely arises due to different perspectives about this type of crime, rather than because female judges are in general more punitive than their male counterparts.
@@
We find evidence of two drivers of the gender conviction rate gap. The first is sharing an identity with the typical victim of domestic violence cases. To investigate if the salience of gender identity matters, we test if the gender difference in strictness is more pronounced in intimate partner violence cases. The incidence and tolerance of this type of offense is strongly associated with patriarchal gender norms involving partners (Heise and Kotsadam, 2015; González and Rodríguez-Planas, 2020). To do this, we take advantage of two key features of our dataset. First, judges are randomly allocated to cases, which makes the relationship between the victim and the assailant independent of judge characteristics. Second, we are able to identify the victim–assailant relationship in cases that result in conviction. We find a statistically significant increase (37.5%)in the likelihood of conviction in a case involving intimate partner violence if the judge is a woman. However, the probability of conviction in a case not involving intimate partner violence is the same for men and women judges.
@@
Second, we test for gender differences in interpretation of the evidence provided in domestic violence cases. To test for the relevance of this driver, we compare the gender conviction rate between cases in which the defendant was caught by the police while committing the crime and cases without this objective piece of evidence. We find that male and female judges make similar decisions when analyzing cases where defendants are caught in the act, while judges differ when making decisions about cases with more subjective evidence. Hence, female judges interpret evidence differently from their male peers when analyzing domestic violence cases. This may mean that judges can make different decisions about what constitutes “true domestic violence” versus victims “gaming the system,” as characterized in Kafka et al. (2019), or that judges may disagree about the appropriate severity of punishment when there is room for discretion.
@@
Finally, we document that the difference in conviction rates between male and female judges is more pronounced later in their careers, which is when career pressures are weaker. We define late-career judges as those in the last stage of the career of a Trial Court judge. Because judges in Brazil are not elected and have job security, [^2] By law, they acquire job stability after 2 years on the job. The goal is to insulate them from political pressure. promotion is the main form of career incentive. However, promotion from the Trial Court to the Appeals Court is rare and extremely unlikely in the short run, unlike career progression at earlier stages. We find that there is no gender conviction rate gap for junior judges, while senior female judges convict domestic violence cases more frequently than male judges in the same career stage. This is an important avenue for future research because it could mean that the difference in perspective of early-career judges is not entirely reflected in the gender conviction gap, potentially because of early-career judges’ desire to conform to the court’s established norms.
@@
This article contributes to the literature about in-group bias in the judicial system. A large body of work finds evidence of judges making different decisions when they share a group identity with defendants or plaintiffs (Kruttschnitt and Savolainen, 2009; Shayo and Zussman, 2011a, 2017; Bielen and Grajzl, 2020; Cai et al., 2022; Chen et al., 2022). Similarly to our work, Kruttschnitt and Savolainen (2009), Bielen and Grajzl (2020), and Chen et al. (2022) analyze in-group gender bias in cases of uncommon crime types (rape and human trafficking), civil cases, and for female offenders, respectively. Our work differs from theirs in that we focus on a single common crime type (domestic violence) whose victims are mostly women.
@@
The paper is organized as follows. Section 2 describes our dataset, while Section 3 explains our empirical strategy. In Section 4, we document the gender conviction rate gap, while 5 discusses its driving forces. Section 7 concludes.
@@
The data also contain two pieces of information that are crucial for the construction of our sample, which we discuss in detail in Section 2.3. The first is a dummy indicating if the case was randomly assigned to a judge within a court district and point in time [^3] That is, not connected to other crime randomly assigned previously.. The second is the name of the courthouse, where different courthouses correspond to certain legal specializations. We use this information to make sure that domestic violence cases are indeed randomly assigned within a district and not systematically assigned to (or reassigned to) judges who do not preside over cases of other types of crimes.
@@
First, we restrict the dataset to cases that were randomly allocated to judges. By Brazilian Law, a case can be allocated to judges by either connected distribution and free distribution. In the first procedure, the crime under consideration is connected to an existing criminal process. Here, the current case is then attached to the previous process and analyzed by the judge associated with the older case. [^8] For two simple examples, a case where a murderer then stole a car to flee the police or a case where a drug trafficker was then caught laundering money. In the second procedure, the crime under consideration is not connected to an existing criminal process. The new case is then allocated to a judge working in the court district where the offense occurred. If there is more than one judge working in that court district, Brazilian Law dictates that the case be randomly allocated to any judge in the court district. To ensure that the cases in our sample are randomly allocated to judges (and therefore uncorrelated with judge gender), we focus on cases that are allocated according to the free distribution system.
@@
Last, we need to ensure that the assigned judge’s gender is determined correctly. To do this, we focus on cases were allocated to judges between January 2011 and December 2019 and use our productivity dataset to remove all observations in which the assigned judge was not active at the court district and month of the assignment. [^9] Cases are randomly allocated to courtrooms and instead of directly to individual judges. If a court district is understaffed at the time of case assignment and there is no presiding judge in the assigned courtroom assigned, the system may record the case as being attributed to the next person who takes over the courtroom, which may take place at a later point in time.
Using this final sample, we compute summary statistics for each possible case outcome and crime type. Table 1, displays case outcomes in the columns and crime types in the rows. Domestic violence cases are more frequently acquitted than other misdemeanors or physical assault cases. As conviction rates are similar across types of crimes, this higher acquittal is due to fewer NPAs and dismissed cases.
Table 1: Case Outcomes by Crime Type in %
Crime Type Conviction NPA Acquittal Dismissal Expired Dead # Cases
(1) (2) (3) (4) (5) (6)
DV 32.5 1.1 51.4 4.3 8.8 2.0 4,358
Other Mis 36.9 5.4 31.2 15.5 9.5 2.2 39,815
Other PA 27.2 7.0 34.1 20.6 10.4 1.9 1,648
  • Notes: The rows of this table indicate the different samples used in this paper: domestic violence cases (DV), misdemeanor cases unrelated to domestic violence (Mis), and physical assault cases that are not domestic violence (PA). The first six columns indicate the frequency of each type of case outcome. The conviction dummy equals one when at least one defendant is convicted in at least one charge. The NPA dummy equals one when the defendant and the district attorney’s office sign an NPA. The dismissal dummy equals one if the case is dismissed. The acquittal dummy equals one if the defendant is found not guilty. The “expired” dummy equals one if at least one of the charges in the case has expired. The “Dead” dummy equals one if the defendant died the sentence’s date. The rows do not sum to 100% because cases with multiple charges can have different outcomes for each charge.
@@
In this section, we explain the empirical strategy used to measure the gender conviction rate gap and indirectly test the strategy’s validity.
@@
Our parameter of interest, $\beta$, measures whether a female judge is more likely to convict than a male judge under the assumption that the judge’s gender is independent of the case’s potential outcome. This exogeneity assumption is valid according to Brazilian Law mandating cases are randomly allocated to judges within court districts (Section 2.3). We indirectly test the validity of this assumption in Table 2.
Even when our exogeneity assumption is valid, the $\beta$ parameter in Equation (1) may capture differences between male and female judges that are unrelated to their views on domestic violence. To measure whether this conviction rate gap is due to different gender perspectives about this type of crime, we compare this domestic violence conviction gap against the gap for similar crime types. Our sample therefore consists of domestic violence cases and one type of comparable criminal cases. We estimate the following linear regression model:
@@
In Equation (2), our parameter of interest is $\beta_{3}$. This specification accounts for the extra strictness and social unacceptability of domestic violence offenses beyond that of other violent crimes of comparable severity. Consequently, this coefficient captures the part of the gender conviction rate gap that is likely due to different gender perspectives about domestic violence cases.
@@
Quarter x Court District FE YES YES YES
@@
  • Notes: Each cell represents the linear relationship between the gender of the judge randomly assigned to the case and a set of characteristics of the case. Column 1 shows this relationship for the sample of all physical assault crimes, Column 2 shows this relationship for the sample of misdemeanors (crimes such that the maximum sentence does not exceed 4 years). Column 3 shows this relationship for the sample of domestic violence crimes. All regressions include quarter-by-court district fixed effects.
Measuring the Gender Conviction Rate Gap
@@
Table 3 shows the effect of judge gender on the probability of conviction in domestic violence cases according to Equation (1). In the first column, we find that randomly selecting a woman or a man to preside does significantly impact conviction rates.
@@
Experience $0.01^{***}$ $0.01^{***}$
$(0.00)$ $(0.00)$
Years in Career Stage $-0.01^{*}$ $-0.01^{*}$
$(0.01)$ $(0.01)$
Caught Red Handed? $0.06$
$(0.04)$
Defendant in jail $-0.03$
$(0.04)$
@@
Quarter x Court District FE YES YES YES
Career Stage Controls NO YES YES
@@
  • Notes: This table presents the estimated results of Equation (1). The first column shows the effect of judge gender on the probability of conviction. The second column includes controls for judge characteristics (career stage, number of years of experience as a judge, and number of years of experience in the current career stage). The third column controls for both judge and case characteristics. All regressions include quarter-by-court district fixed effects.
@@
Career Controls NO YES YES NO YES YES
Case Controls NO NO YES NO NO YES
Quarter-District FE YES YES YES YES YES YES
@@
The interaction coefficient captures the part of the gender conviction rate gap that is specific to domestic violence cases and is likely due to different approaches to domestic violence cases between male and female judges. In the next section, we explore this difference more deeply.
@@
In this section, we show two important findings about the conviction rate gender gap. The first is suggestive evidence that the gap is explained by judges’ identification with their own gender. We find that the difference in conviction rates between male and female judges is more intense in domestic violence cases involving intimate partner violence. Violence between relationship partners is highly impacted by patriarchal social norms that make gender identity more salient.
The second finding, discussed in Section 5.2, is that men and women respond differently to evidence presented in the same case. We present evidence that female judges are more likely than male judges to convict in cases where evidence is more tenuous and subject to interpretation, but not when the evidence is objectively incriminating.
@@
Next, we show that the gender of the judge matters more in cases where patriarchal social norms are particularly relevant: intimate partner violence. We interpret this finding as indicative that a major reason for the gender conviction gap is that domestic violence cases make gender identity more salient to judges.
@@
First, sexist social norms and gender roles significantly affect the occurrence and punishment of intimate partner violence in multiple contexts. For instance, González and Rodríguez-Planas (2020) show that, among immigrant families, the incidence of intimate partner violence is positively affected by gender norms in the country of origin. Heise and Kotsadam (2015) show that the incidence of intimate partner violence in different countries is strongly associated with patriarchal social norms, such as male authority over women, justifications for wife beating, and laws and practices that disadvantage women in property ownership and other economic resources. Closer to our context, Perova and Reynolds (2017) show that police stations that specialize in domestic violence with specially trained officers significantly reduce the number of deaths of young women by intimate partner violence in metropolitan municipalities in Brazil. In terms of judicial decision-making on domestic violence, the negative stereotypes referred to in Kafka et al. (2019) and HumanRightsWatch (1991) concern female intimate partners.
@@
We then take advantage of an important feature of our data: we are able to identify the relationship between the victim and the assailant for all cases that culminate in conviction. [^10] In fact, we observe this relationship for cases in which the judge must make an active decision, e.g., excluding cases where the judge declares the defendant dead, a non-prosecution agreement is reached, or the case expires. To do this, we construct an algorithm that defines a case as intimate partner violence if, when writing the court’s decision, the judge uses the Portuguese word for wife, girlfriend, ex-wife, ex-girlfriend, or similar words. [^11] In the domestic violence sample, 57.8% of cases are flagged as violence against partners or ex-partners. The specific words we look for are “namorada,” “ex-namorada,” “esposa,” “ex-esposa,” “amásia,” “ex-amásia,” “companheira,” “ex-companheira,” “mulher,”, “ex-mulher,” “coabitação,” “união estável,” “casamento,” “era casado,” and “era casada.”
To show that the gender of the judge matters more for the outcome of intimate partner violence cases than for that of other domestic violence offenses, we make the following argument. If the probability of a conviction for non-intimate partner violence is non-zero, then two things cannot be true at the same time: 1) the probability of conviction for a domestic violence offense involving intimate partner violence is higher for women than it is for men; 2) the probability of conviction for a domestic violence offense not involving an intimate partner is not higher for women than it is for men. Although we cannot observe the relationship between victim and assailant in all cases, we can reasonably assume that this relationship is independent of the assigned judge’s gender due to the fact that cases are randomly allocated to judges in our sample.
@@
To test for this possibility, we interact the indicator for assault between intimate partners with the conviction indicator. This interaction term becomes our new outcome variable in Equation (1).
@@
Quarter-District FE YES YES YES
Case Controls YES YES YES
Judge Career Controls YES YES YES
@@
These findings are consistent with existing evidence that, in many contexts, judges make different decisions when they share a group identity with defendants or plaintiffs (Kruttschnitt and Savolainen, 2009; Shayo and Zussman, 2011a, 2017; Bielen and Grajzl, 2020; Cai et al., 2022; Chen et al., 2022). Many of these papers show that the differences intensify as group identity becomes more salient. This intensification may result from of a heightened opposition between two or more social groups (“us” versus “them”). For example, Shayo and Zussman (2011b) show that Jewish Israeli judges in small claim courts favor Jewish litigants more intensely when they are more exposed to conflict-related terrorism. Closer to our case, Cai et al. (2022) show that judges analyzing divorce cases in China favor plaintiffs of their own gender more in areas where social norms are more conservative. In our case, we show that the gender gap is much more intense in cases where the form of domestic violence is more stereotypical [^12] Note that most domestic violence against women is committed by partners and ex-partners (United Nations Office on Drugs and Crime, 2023). and has the potential to cause the most division because of social norms. Judges’ views on other forms of domestic violence may be less dependent on their views on patriarchy. For example, beating one’s own parents or children is more likely to be viewed as unjustifiable by men and women alike.

### v2

@@ Abstract
We investigate the impact of judges’ gender on the outcome of domestic violence cases. Using data from São Paulo, Brazil, between 2010 and 2019, we compare conviction rates by judge’s gender and find that a domestic violence case assigned to a female judge is 31% (10 p.p) more likely to result in conviction than a case assigned to a male judge with similar career characteristics. To show that this decision gap rises due to different gender perspectives about domestic violence instead of rising due to female judges being tougher than male judges, we compare it against gender conviction rate gaps in similar types of crimes. We find that the gender conviction rate gap for domestic violence cases is significantly larger than the same gap for other misdemeanor cases (3 p.p. larger) and for other physical assault cases (8 p.p. larger). Lastly, we find evidence that at least two channels explain this gender conviction rate gap for domestic violence cases: gender-based differences in evidence interpretation and gender-based sentencing criteria.
@@ Introduction
Violence against women is frequently perpetrated inside the home by the ones closest to them. The United Nations Office on Drugs and Crime (2023) estimates that 55% of all homicides against women worldwide in 2022 were committed by intimate partners or other family members. As this violence is a criminal offense, key components of the fight against domestic violence are the Criminal Court System and how court judges approach this type of crime (Beecher-Monas, 2001). For instance, the Human Rights Watch (1991) argues that, in the 1980s and early 1990s, Brazilian judges failed to analyze crimes of domestic violence against women in a non-discriminatory manner. [^1] In the years since, Brazilian authorities have made important legal and institutional changes to address domestic violence crimes. The most prominent of these is the Maria da Penha Law, enacted in 2006, which characterizes domestic violence as a specific type of criminal offense. Moreover, Kafka et al. (2019) finds that American judges in North Carolina use gender stereotypes to argue that some women want to “game the system” with “frivolous cases” when requesting protective orders against their abusive partners.
@@
In this analysis, we find that a domestic violence case assigned to a female judge is 31% (10 p.p) more likely to result in conviction than a case assigned to a male judge with similar career characteristics. Since this decision gap may be due to differences between male and female judges regardless of their perspectives on domestic violence, we compare this gap to gender decision gaps for two other types of crime: misdemeanors and physical assault. Because we focus on domestic violence crimes that do not result in death or long-term physical impairment, misdemeanors are comparable to domestic violence cases in their sentence severity (according to Brazilian Law). Physical assault cases are comparable to domestic violence cases because both involve some sort of physical aggression.
We find that the gender conviction-rate gap for domestic violence cases is more than twice the size of that of other misdemeanors. Furthermore, while there is a gender conviction-rate gap for domestic violence cases, there are no differences between female and male judges when analyzing other types of physical assault cases. Consequently, the gender conviction-rate gap for domestic violence cases likely arises due to different perspectives about this type of crime rather than because female judges are, in general, more punitive than their male counterparts.
For this reason, we classify this type of gap as an in-group bias according to the definition given by Shayo and Zussman (2011). We do so because victims of domestic violence are overwhelmingly female, and members of this group (female judges) act differently from members of the other group (male judges) when analyzing this type of crime specifically.
We then investigate two possible drivers of this in-group bias. They were proposed by Boyd, Epstein, and Martin (2010) and are known as the representational account and the informational account. The first one argues that female judges act as representatives of their group and tend to protect it more intensely than male judges. The second one argues that female judges process the information contained in domestic violence cases in a unique way. Our evidence suggests that both accounts drive the in-group bias in domestic violence cases in São Paulo, Brazil.
@@
To investigate if the salience of gender identity matters (representational account), we test if the gender difference in strictness is more pronounced in intimate partner violence cases. To do so, we explore each case’s full sentence and estimate whether female judges are more likely to convict a defendant for a domestic violence offense and simultaneously flag the case as associated with intimate partner violence. We find that the gender conviction rate gap is 37.5% higher (9 p.p.) in cases in which a female judge may identify more strongly with the victim than a male judge, while there is no gender difference in cases that are not flagged as intimate partner violence offenses. This difference is associated with the representational account of the in-group bias because the incidence and tolerance of this type of offense are strongly associated with patriarchal gender norms involving partners (Heise and Kotsadam, 2015; González and Rodríguez-Planas, 2020).
@@
To test for gender differences in the interpretation of the evidence provided in domestic violence cases (informational account), we compare the gender conviction rate between cases in which the defendant was caught by the police while committing the crime and cases without this objective piece of evidence. We find that male and female judges make similar decisions when analyzing cases where defendants are caught in the act, while they differ when making decisions about cases with more subjective evidence. Hence, female judges process the information contained in domestic violence cases differently from their male peers, as suggested by the informational account of the in-group bias.
@@
Finally, we document that the difference in conviction rates between male and female judges is more pronounced later in their careers, which is when career pressures are weaker. We define late-career judges as those in the last stage of the career of a Trial Court judge. Because judges in Brazil are not elected and have job security, [^2] By law, they acquire job stability after 2 years on the job. The goal is to insulate them from political pressure. promotion is the main form of career incentive. However, promotion from the Trial Court to the Appeals Court is rare and extremely unlikely in the short run, unlike career progression at earlier stages (Laneuville, 2024). We find that there is no gender conviction-rate gap for junior judges, while senior female judges convict domestic violence cases more frequently than male judges in the same career stage. This is an important avenue for future research because it could mean that the difference in perspective of early-career judges is not entirely reflected in the gender conviction-rate gap, potentially because of early-career judges’ desire to conform to the court’s established norms.
@@
This article contributes to the literature on judicial biases. Harris and Sen (2019) reviews this literature, focusing on ideology, race and gender gaps. Within the studies focusing on in-group bias, many authors find evidence of judges making different decisions when they share a group identity with defendants or plaintiffs (Kruttschnitt and Savolainen, 2009; Boyd, Epstein, and Martin, 2010; Shayo and Zussman, 2011, 2017; Bielen and Grajzl, 2020; Cai et al., 2022; Chen et al., 2022). Similarly to our work, Kruttschnitt and Savolainen (2009), Boyd, Epstein, and Martin (2010) Bielen and Grajzl (2020), and Chen et al. (2022) analyze in-group gender bias in cases of uncommon crime types (rape and human trafficking) in Finland, federal appellate cases in the U.S., civil cases in Belgium, and for female offenders in Kenya, respectively. Our work differs from theirs in that we focus on a single common crime type (domestic violence) whose victims are mostly women. Furthermore, while Boyd, Epstein, and Martin (2010) finds that in-group bias in federal appellate cases in the U.S. is consistent with the informational account only, we find that in-group bias in criminal cases in Brazil is consistent with the representational and informational accounts.
@@
The paper is organized as follows. Section 2 describes our dataset, while Section 3 explains our empirical strategy. Moreover, Section 4 documents the gender conviction rate gap, while Section 5 discusses its driving forces. Finally, Section 6 argues that the difference in perspective of early-career judges may not be entirely reflected in the gender conviction-rate gap, and Section 7 concludes.
@@
The data also contain two pieces of information that are crucial for the construction of our sample, which we discuss in detail in Section 2.3. The first is a dummy indicating if the case was randomly assigned to a judge within a court district and point in time. [^3] Alternatively, some cases are connected to other crimes and are assigned to the judge who was assigned to the previous criminal case. The second is the name of the courtroom, where different courtrooms may correspond to certain legal specializations. We use this information to make sure that domestic violence cases are indeed randomly assigned within a district and not systematically assigned to (or reassigned to) judges who do not preside over cases of other types of crimes.
@@
First, we restrict the dataset to cases that were randomly allocated to judges. By Brazilian Law, a case can be allocated to judges by either connected distribution or free distribution. In the first procedure, the crime under consideration is connected to an existing criminal process. Here, the current case is then attached to the previous process and analyzed by the judge associated with the older case. [^8] For two simple examples, consider a case where a murderer killed someone and then stole a car to flee the police or a case where a drug trafficker was also caught laundering money. In the second procedure, the crime under consideration is not connected to an existing criminal process. The new case is then allocated to a judge working in the court district where the offense occurred. If there is more than one judge working in that court district, Brazilian Law dictates that the case be randomly allocated to any judge in the court district. To ensure that the cases in our sample are randomly allocated to judges (and therefore uncorrelated with judge gender), we focus on cases that are allocated according to the free distribution system.
@@
Last, we need to ensure that the assigned judge’s gender is determined correctly. To do this, we focus on cases allocated to judges between January 2011 and December 2019 and use our productivity dataset to remove all observations in which the assigned judge was not active at the court district and month of the assignment. [^9] Cases are randomly allocated to courtrooms and instead of directly to individual judges. If a court district is understaffed at the time of case assignment and there is no presiding judge in the assigned courtroom, the system may record the case as being attributed to the next person who takes over the courtroom, which may take place at a later point in time.
Using this final sample, we compute summary statistics for each possible case outcome and crime type. Table 1 displays case outcomes in the columns and crime types in the rows. It also reports the share of female judges analyzing each type of crime. Domestic violence cases are more frequently acquitted than other misdemeanors or physical assault cases. As conviction rates are similar across types of crimes, this higher acquittal is due to fewer NPAs and dismissed cases. We also note that the share of female judges is similar across crime types and that there are 270 female judges and 425 male judges in our entire sample.
Table 1: Case Outcomes and Share of Female Judges by Crime Type in %
Crime Conviction NPA Acquittal Dismissal Expired Dead Female Obs
Type Judge
DV 32.5 1.1 51.4 4.3 8.8 2.0 32.1 4,358
Other Mis 36.9 5.4 31.2 15.5 9.5 2.2 36.4 39,815
Other PA 27.2 7.0 34.1 20.6 10.4 1.9 38.1 1,648
  • Notes: The rows of this table indicate the different samples used in this paper: domestic violence cases (DV), misdemeanor cases unrelated to domestic violence (Mis), and physical assault cases that are not domestic violence (PA). The first six columns indicate the frequency of each type of case outcome. The conviction dummy equals one when at least one defendant is convicted in at least one charge. The NPA dummy equals one when the defendant and the district attorney’s office sign an NPA. The dismissal dummy equals one if the case is dismissed. The acquittal dummy equals one if the defendant is found not guilty. The “expired” dummy equals one if at least one of the charges in the case has expired. The “Dead” dummy equals one if the defendant died before the sentence’s date. The rows do not sum to 100% because cases with multiple charges can have different outcomes for each charge. The second to last column shows the share of female judges in each sample while the last column shows the size of each sample.
@@
In this section, we explain the empirical strategy used to measure the gender conviction-rate gap and indirectly test the strategy’s validity.
@@
Our parameter of interest, $\beta$, measures whether a female judge is more likely to convict than a male judge under the assumption that the judge’s gender is independent of the case’s potential outcome. This exogeneity assumption is valid according to Brazilian Law, which mandates cases to be randomly allocated to judges within court districts (Section 2.3). We indirectly test the validity of this assumption in Table 2.
Even when our exogeneity assumption is valid, the $\beta$ parameter in Equation (1) may capture differences between male and female judges that are unrelated to their views on domestic violence. To measure whether this conviction-rate gap is due to different gender perspectives about this type of crime, we compare this domestic violence conviction gap against the gap for similar crime types. Now, our sample consists of domestic violence cases and one type of comparable criminal case. We estimate the following linear regression model:
@@
In Equation (2), our parameter of interest is $\beta_{3}$. This specification accounts for the extra strictness and social unacceptability of domestic violence offenses beyond that of crimes of comparable severity (coefficient $\beta_{2}$) and for the possibility that female judges are stricter than male judges in general (coefficient $\beta_{1}$). Consequently, this coefficient captures the part of the gender conviction-rate gap that is likely due to different gender perspectives about domestic violence cases specifically. For this reason, we interpret the $\beta_{3}$ coefficient as capturing a type of in-group bias according to the definition given by Shayo and Zussman (2011). We do so because victims of domestic violence are overwhelmingly female, and members of this group (female judges) may act differently from members of the other group (male judges) when analyzing this type of crime in particular.
@@
Quarter x Court District FE $\checkmark$ $\checkmark$ $\checkmark$
@@
  • Notes: Each cell represents the linear relationship between the gender of the judge randomly assigned to the case and a set of characteristics of the case. Column (1) shows this relationship for the sample of all physical assault crimes, Column (2) shows this relationship for the sample of misdemeanors (crimes such that the maximum sentence does not exceed 4 years), and Column (3) shows this relationship for the sample of domestic violence crimes. All regressions include quarter-by-court district fixed effects.
Measuring the Gender Conviction-Rate Gap
@@
Table 3 shows the effect of judge gender on the probability of conviction in domestic violence cases according to Equation (1). In the first column, we find that randomly selecting a woman or a man to preside does not significantly impact conviction rates.
@@
Quarter x Court District FE $\checkmark$ $\checkmark$ $\checkmark$
Career Controls $\checkmark$ $\checkmark$
Case Controls $\checkmark$
@@
  • Notes: This table presents the estimated results of Equation (1). The first column shows the effect of judge gender on the probability of conviction. The second column includes controls for judge characteristics (career stage, number of years of experience as a judge, and number of years of experience in the current career stage). The third column controls for both judge and case characteristics (whether the defendant was caught red-handed and whether the defendant was arrested by the sheriff before the trial). All regressions include quarter-by-court district fixed effects.
@@
Quarter-District FE $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Career Controls $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Case Controls $\checkmark$ $\checkmark$
@@
The interaction coefficient captures the part of the gender conviction rate gap that is specific to domestic violence cases and is likely due to different approaches to domestic violence cases between male and female judges. For this reason, this result suggests the existence of a type of in-group bias according to the definition given by Shayo and Zussman (2011). [^10] In Appendix C, we propose a simple economic model formalizing the definition of in-group bias. In the next section, we explore the driving forces behind this in-group bias.
@@
In this section, we show two important findings about the conviction rate gender gap and its driving forces.
The first is suggestive evidence that the gap is explained by judges’ identification with their own gender (Section 5.1). We find that female judges are more likely than male judges to convict and simultaneously flag the case as an intimate partner violence offense, while female and male judges are equally likely to convict and not flag the case as an intimate partner violence offense. This result suggests that the representational account (Boyd, Epstein, and Martin, 2010) is one of the explanatory forces of the in-group bias documented in Table 4 because violence between relationship partners is highly impacted by patriarchal social norms that make gender identity more salient (Heise and Kotsadam, 2015; González and Rodríguez-Planas, 2020).
The second finding, discussed in Section 5.2, is that men and women respond differently to evidence presented in the same case. We present evidence that female judges are more likely than male judges to convict in cases where evidence is more tenuous and subject to interpretation, but not when the evidence is objectively incriminating. This result suggests that the informational account (Boyd, Epstein, and Martin, 2010) is one of the explanatory forces of the in-group bias documented in Table 4.
In Appendix C, we propose a simple threshold-crossing model that formalizes the concepts of the representational and informational accounts.
@@
In this section, we show that the gender of the judge matters more in cases where patriarchal social norms are particularly relevant: intimate partner violence. We interpret this finding as indicative that a major reason for the gender conviction gap is that domestic violence cases make gender identity more salient to judges, suggesting that the representational account (Boyd, Epstein, and Martin, 2010) is one of the explanatory forces of the in-group bias documented in Section 4.
@@
We argue that domestic violence cases make gender identity more salient because sexist social norms and gender roles significantly affect the occurrence and punishment of intimate partner violence in multiple contexts. For instance, González and Rodríguez-Planas (2020) show that, among immigrant families, the incidence of intimate partner violence is positively affected by gender norms in the country of origin. Heise and Kotsadam (2015) show that the incidence of intimate partner violence in different countries is strongly associated with patriarchal social norms, such as male authority over women, justifications for wife beating, and laws and practices that disadvantage women in property ownership and other economic resources. Closer to our context, Perova and Reynolds (2017) show that police stations that specialize in domestic violence with specially trained officers significantly reduce the number of deaths of young women by intimate partner violence in metropolitan municipalities in Brazil. In terms of judicial decision-making on domestic violence, the negative stereotypes mentioned by Kafka et al. (2019) and Human Rights Watch (1991) concern female intimate partners.
@@
To show that the gender of the judge matters more for the outcome of intimate partner violence cases than for that of other domestic violence offenses, we take advantage of an important feature of our data. We are able to identify the relationship between the victim and the assailant for all cases that culminate in an active decision by the presiding judge (conviction or acquittal). To do this, we construct an algorithm that defines a case as intimate partner violence if, when writing the court’s decision, the judge uses the Portuguese word for wife, girlfriend, ex-wife, ex-girlfriend, or similar words. [^11] In the domestic violence sample, 57.8% of cases are flagged as violence against partners or ex-partners. The specific words we look for are “namorada,” “ex-namorada,” “esposa,” “ex-esposa,” “amásia,” “ex-amásia,” “companheira,” “ex-companheira,” “mulher,”, “ex-mulher,” “coabitação,” “união estável,” “casamento,” “era casado,” and “era casada.”
We, then, test whether female judges are more likely to make a conviction decision in a domestic violence case and flag the case as being related to intimate partner violence after controlling for case characteristics. If there is a gender conviction gap in cases flagged as intimate partner violence but not in cases without such a flag, then a stronger identification with the victim may be driving the gender lenience gap found in Section 4. If the gender conviction rate gap is present regardless of explicit mentions of intimate partner violence, then there must be other forces driving this lenience gap.
@@
To implement this test, we interact the indicator for assault between intimate partners with the conviction indicator. This interaction term becomes our new outcome variable in Equation (1).
@@
Quarter-District FE $\checkmark$ $\checkmark$ $\checkmark$
Case Controls $\checkmark$ $\checkmark$ $\checkmark$
Judge Career Controls $\checkmark$ $\checkmark$ $\checkmark$
@@
These findings are consistent with existing evidence that, in many contexts, judges make different decisions when they share a group identity with defendants or plaintiffs (Kruttschnitt and Savolainen, 2009; Shayo and Zussman, 2011, 2017; Bielen and Grajzl, 2020; Cai et al., 2022; Chen et al., 2022). Many of these papers show that the differences intensify as group identity becomes more salient. This intensification may result from of a heightened opposition between two or more social groups (“us” versus “them”). For example, Shayo and Zussman (2011) show that Jewish Israeli judges in small claim courts favor Jewish litigants more intensely when they are more exposed to conflict-related terrorism. Closer to our case, Cai et al. (2022) show that judges analyzing divorce cases in China favor plaintiffs of their own gender more in areas where social norms are more conservative. In our case, we show that the gender gap is much more intense in cases where the form of domestic violence is more stereotypical. [^12] Note that most domestic violence cases against women are committed by partners and ex-partners (United Nations Office on Drugs and Crime, 2023). Consequently, the representational account (Boyd, Epstein, and Martin, 2010) is a plausible driving forcing of the in-group bias documented in Section 4.
@@
In this section, we show that female judges are more likely than male judges to convict in domestic violence cases where evidence is more tenuous and subject to interpretation, but not when the evidence is objectively incriminating. We interpret this result as suggesting that the informational account (Boyd, Epstein, and Martin, 2010) is one of the explanatory forces of the in-group bias documented in Table 4.